OpenStudy (anonymous):

What is an equation of the line, in point-slope form, that passes through the given point and has the given slope? point: (8, –8); slope: 3 y – 8 = 3 (x + 8) y + 8 = 3 (x + 8) y – 8 = 3 (x – 8) y + 8 = 3 (x – 8) I got B am I right?

OpenStudy (noseboy908):

no, fill in point-slope form

OpenStudy (noseboy908):

again, that is

OpenStudy (anonymous):

y – y1 = m(x – x1) (8, –8); slope: 3 y - (-8) = 3(x - 8)

OpenStudy (noseboy908):

yep. now make the two "minuses" into one positive

OpenStudy (anonymous):

y + 8 = 3(x - 8) So D?

OpenStudy (noseboy908):

yep!